OpenStudy (anonymous):

In 1920, the record for a certain race was 45.4 seconds. In 1980, it was 44.2 seconds. Let R(t) = the record in the race and t=the number of years since 1920. Find the linear function that fits the data: R(t) = ? Record for 2003? Record for 2006? And in what year will the predicted record b e 43.64 seconds?

OpenStudy (a_clan):

R(t)= 45.4 - 0.02t for 2003, t= (2003-1920)=83 So, R(83) = 45.4 - 0.02*83 = 43.74 Similarly for 2006, t= 86